OpenStudy (anonymous):

How to solve this: 1/2(4)^2 - 1/3(4)^3 +12(4) - 1/2(-3)^2 +1/3(-3)^3-12(-3)

OpenStudy (anonymous):

Please see the attached file for neat view.

OpenStudy (shubhamsrg):

simple +tion -traction..wont take more than 10-20 secs ??

OpenStudy (shubhamsrg):

whats the catch ?

OpenStudy (anonymous):

subhamsrg the catch is to solve this in the best way possible. There are a few ways to solve this question. you can just calculate term by term and adding it all up in the end. Or, smarter, to factorise and simplying the equation, making the calculate a lot less tedious. This is better sometimes as it lowers your chances of making careless mistakes.

OpenStudy (anonymous):

I don't have idea how to multiply and division apply here. Like 1/2(4)^2 .. in this what we do first solve exponential 4^2 = 16 .. then multiply 16 with 2 .. right ?

OpenStudy (anonymous):

You solve the brackets involving the exponents first, so it looks like this: 1/2(16)-1/3(64)+12(4)-1/2(9)+1/3(9)-12(-3) Now we must do the multiplication first following BEDMAS, so it would look like this: 16/2-64/3+48-9/2+9/3+36 Now we just do the addition/subtraction: 8-64/3+48-9/2+9/3+36 64/3-9/2+9/3+92 Make denominators the same: 384/18-81/18+54/18+1656/18 = 2013/18 That's your answer.

OpenStudy (shubhamsrg):

well if you looking for shortcuts,,you may collect like terms : i.e take 1/2 common from 2 terms 1/3 from other 2 and 2 from remaining 2 you'll have a similpified form.. i'd go for orthodox method as it saves time for me

OpenStudy (anonymous):

I got this answer is this correct 57.167 ?

OpenStudy (shubhamsrg):

correct..

OpenStudy (anonymous):

Thank you all of you for your help, I am confused which one is the best answer so please don't mind :)